evaluate the following integral
∫ cos x / 2 + 3 sin x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{cosx}{2+sinx}dx
2+sinx = t
cosxdx = dt
I = \int \frac{1}{t}dt
I = log(t) + c
I = log(sinx+2) + c